jasonmitchell:

The coordinates of the vertices of △JKL are J(−5, −1) , K(0, 1) , and L(2, −5) . Which statement correctly describes whether △JKL is a right triangle? A. △JKL is a right triangle because JL¯¯¯¯¯ is perpendicular to KL¯¯¯¯¯ . B. △JKL is not a right triangle because no two of its sides are perpendicular. C. △JKL is a right triangle because JK¯¯¯¯¯ is perpendicular to JL¯¯¯¯¯ . D. △JKL is a right triangle because JK¯¯¯¯¯ is perpendicular to KL¯¯¯¯¯ .

jasonmitchell:

@563blackghost

Cupcake123456:

To find out if they form a right triangle, find out if two of the sides are perpendicular. to find out if two sides are perpendicular, find the slopes, if the slopes are negative reciprocals, the two lines are perpendicular. slope of JK: 2/5 slope of KL: -6/2 slope of JL: -4/7 no negative reciprocal pair, so the triangle is not a right triangle.

563blackghost:

So you want to find the slopes of JK, KL, and LJ. So let's first find JK. You would plug it into slope formula. \(\Large\bf{slope=\frac{1+1}{0+5}}\) what does that equal @jasonmitchell ?

Cupcake123456:

@563blackghost u agree

jasonmitchell:

1 + 1 = 2 0 + 5 = 5

563blackghost:

i do, but i think its best to work it with them in case they dont understand et. @Cupcake123456

Cupcake123456:

kk.

563blackghost:

correct. So JK has a slope of 2/5. Now lets find KL. \(\Large\bf{slope=\frac{-5-1}{2-0}}\)

Cupcake123456:

If we were to graph the points on a graph then the coordinates (0,1) to the point (2,-5) would form a right triangle @563blackghost right ?

jasonmitchell:

-6 and 2

563blackghost:

Correct but now you need to simplify. \(\large\bf{\frac{-6}{2} = -3}\). right? @jasonmitchell

jasonmitchell:

correct

563blackghost:

Okie cx Now we find LJ. \(\large\bf{slope=\frac{-5+1}{2+5}}\) what does that equal?

jasonmitchell:

-4 and 7

563blackghost:

Correct. So in order for a triangle to be right is that you have to identify its reciprocal. You flip the fraction and change the signs. So you have `JK = 2/5` `KL = -3` and `LJ = -4/7` identify its reciprocals. For the first one you would flip the fraction then the sign. \(\Large\bf{\frac{2}{5} ~~ \rightarrow ~~ \frac{5}{2} ~~\rightarrow ~~ \color{red}{-}\frac{5}{2}}\) and follow through with each... \(\large\bf{-3 \rightarrow \frac{1}{3}}\) \(\large\bf{-\frac{4}{7} \rightarrow \frac{7}{4}}\) None of these slopes are equal to the slopes of the triangle meaning that this triangle is not a right triangle.

jasonmitchell:

B

563blackghost:

Correct.

563blackghost:

Graphing does work, but you cant assume a right triangle from a visual, you need to find the slopes in order to find if it is a true perpendicular @Cupcake123456

jasonmitchell:

The coordinates of the vertices of quadrilateral JKLM are J(−3, 2) , K(3, 5) , L(9, −1) , and M(2, −3) . Which statement correctly describes whether quadrilateral JKLM is a rhombus? A. Quadrilateral JKLM is not a rhombus because there are no pairs of parallel sides. B. Quadrilateral JKLM is a rhombus because opposite sides are parallel and all four sides have the same length. C. Quadrilateral JKLM is not a rhombus because there is only one pair of opposite sides that are parallel. D. Quadrilateral JKLM is not a rhombus because opposite sides are parallel but the four sides do not all have the same length.

jasonmitchell:

and right i agree

563blackghost:

oh and pls do credit your sources @Cupcake123456 https://brainly.com/question/6753445

563blackghost:

can you post this question in a new thread @jasonmitchell ?